# PQMass: Probabilistic Assessment of the Quality of Generative Models using Probability Mass Estimation

### v2

@@ Appendix G PQMass in different metrics
The PQMass test is performed using only distances between sample points, thus one may choose any distance metric, or kernel (representation) function, to perform the test. The choice of metric may affect what features PQMass is most sensitive to. In Fig. 20 we show how PQMass performs on a toy problem using various distance metrics. A Gaussian mixture model with 20 components is created in 100 dimensions, for each component the means are drawn from a uniform distribution with range $(-10,10)$ in all dimensions, the covariance matrices are constructed with eigenvectors aligned randomly and eigenvalues drawn from a log-uniform distribution with range $(-1,1)$. The weight for each component is also drawn from a log-uniform distribution with range $(-1,1)$. For each test, we draw two sets of $2^{12}$ samples from the GMM as input to PQMass. The null test takes these inputs as-is, for the “scale” out-of-distribution test the samples are scaled by 1.08, and for the “rotated” out-of-distribution test the samples are rotated by approximately 0.08 radians on a random axis (scale and rotation chosen arbitrarily for figure quality). To generate the histograms in the figure, the tests are re-run $2^{14}$ times each.
All metrics in Fig. 20 correctly fail to reject the null test (left subplot), as should be expected.
@@
We also repeat the experiment described in § 3.3, where we evaluate the impact of various distance metrics on PQMass as Gaussian noise with increasing variance is added to the ImageNet dataset. In Figure  Fig. 21, we show that when no noise is added (null test), all distance metrics will correctly detect that the images are in distribution (left subplot). We also show that as we add noise, different metrics showcase different sensitivity and thus affect $\chi_{\rm PQM}^{2}$ differently. In both tests (and indeed all our tests) the Euclidean distance – the one used in the main experiments – is among the most discriminating.
@@ Appendix H Comparison of PQMass to other two-sample tests
Figure 22: Comparison of 50 samples from $\mathcal{N}(\mathbf{0},\mathbf{I})$ to 50 samples from $\mathcal{N}(\mathbf{0.5},\mathbf{I})$, with PQMass. Unlike the multivariate two-sample test shown in Table 6, PQMass can confidently detect that the two sets of samples are indeed not from the same distribution, showcasing that PQMass is more sensitive than the multivariate two-sample tests presented in the main text.
[…]
@@ Appendix I Permutation tests
Throughout our results, a common test metric was the mean of PQMass run with many retessellations, though with the same input data. In principle, these are not independent tests and so should not produce a $\chi^{2}$ distribution with mean equal to $n_{r}-1$. In practice, we find that for high-dimensional problems, the retessellations without resampling very closely approximate a $\chi^{2}$ distribution (in the null case), though this is not the case in low dimensions. However, for a more robust test using retessellations without resampling, one can perform permutations of the input samples to determine a valid range for the mean-of-retessellations measurement which is demonstrated in Fig. 23.
In this permutation test, one first computes the mean-of-retessellations like normal (“test stat” in the figure), then for $n_{permute}$ iterations one shuffles the input samples randomly and reruns the mean-of-retessellations measurement (creating the “permute stats” histogram in the figure). The permutation tests satisfy the null hypothesis by construction since they are a mixture of the two inputs.
Under the null hypothesis, the “test stat” must be uniformly sampled from the “permute stats” since all the samples from both inputs are independent and identically distributed. This can concretely determine a p-value based on the mean-of-retessellations. However, this removes the information from the shape of the distribution (one could perform a similar test with std-of-retessellations) and so it is still advised to look at the $\chi^{2}_{\rm PQM}$ distribution for irregularities. Further, this requires $n_{permute}$ times the compute to complete the test, which can become very large in real scenarios. Ultimately, we find PQMass is sensitive enough that in most cases the “permute test” is unnecessary and the rejection of the null hypothesis can be conservatively done based on the $\chi^{2}(n_{r}-1)$ distribution.
@@
Figure 23: Permutation test for validity of taking the mean of many PQMass retessellations. Samples are drawn from a 100 dimension standard Gaussian, we use 1000 samples for each input. Shown are the $\chi^{2}(99)$ distribution (black dashed), mean of many retessellations (red vertical line), and the permutation test (purple histogram). Left: A null test which, as expected, fails to reject the null hypothesis. Right: An out of distribution test where one sample has been scaled by 1.1. The null hypothesis is indeed rejected.

### v3

@@ Appendix G PQMass in different metrics
The PQMass test is performed using only distances between sample points, thus one may choose any distance metric, or kernel (representation) function, to perform the test. The choice of metric may affect what features PQMass is most sensitive to. In Fig.˜20 we show how PQMass performs on a toy problem using various distance metrics. A Gaussian mixture model with 20 components is created in 100 dimensions, for each component the means are drawn from a uniform distribution with range (10,10)(-10,10) in all dimensions, the covariance matrices are constructed with eigenvectors aligned randomly and eigenvalues drawn from a log-uniform distribution with range (1,1)(-1,1). The weight for each component is also drawn from a log-uniform distribution with range (1,1)(-1,1). For each test, we draw two sets of $2^{12}$ samples from the GMM as input to PQMass. The null test takes these inputs as-is, for the “scale” out-of-distribution test the samples are scaled by 1.08, and for the “rotated” out-of-distribution test the samples are rotated by approximately 0.08 radians on a random axis (scale and rotation chosen arbitrarily for figure quality). To generate the histograms in the figure, the tests are re-run $2^{14}$ times each.
All metrics in Fig.˜20 correctly fail to reject the null test (left subplot), as should be expected.
@@
We also repeat the experiment described in §˜3.3, where we evaluate the impact of various distance metrics on PQMass as Gaussian noise with increasing variance is added to the ImageNet dataset. In Figure  Fig.˜21, we show that when no noise is added (null test), all distance metrics will correctly detect that the images are in distribution (left subplot). We also show that as we add noise, different metrics showcase different sensitivity and thus affect $\chi_{\rm PQM}^{2}$ differently. In both tests (and indeed all our tests) the Euclidean distance – the one used in the main experiments – is among the most discriminating.
@@ Appendix H Comparison of PQMass to other two-sample tests
Figure 22: Comparison of 50 samples from $\mathcal{N}(\mathbf{0},\mathbf{I})$ to 50 samples from $\mathcal{N}(\mathbf{0.5},\mathbf{I})$, with PQMass. Unlike the multivariate two-sample test shown in Table˜6, PQMass can confidently detect that the two sets of samples are indeed not from the same distribution, showcasing that PQMass is more sensitive than the multivariate two-sample tests presented in the main text.
We show that PQMass can outperform this methodology by comparing the performance of the unweighted and weighted nearest-neighbor two-sample tests against PQMass. Specifically, we consider two distributions, $\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\mathcal{N}(\mathbf{0.5},\mathbf{I})$, where $\mathbf{0}=(0,0)$ and $\mathbf{0.5}=(0.5,0.5)$ are 2-dimensional mean vectors and $\mathbf{I}$ denotes the $2x2$ identity matrix. Samples from these distributions should be identified as out-of-distribution relative to each other. Using 50 samples from each distribution, we evaluate $T_{k,n}$ and $U_{k,n,w}$ in Table˜6. For the unweighted test statistics, the expected value is $T_{k,n}$ = 49.4900, and the observed value is $T_{k,n}$ = 52.4000, with a p-value of $\text{p-value}_{T_{k,n}}=0.5596$. These results indicate that the unweighted two-sample test fails to reject the null hypothesis and incorrectly suggests the distributions are in-distribution. For the weighted two-sample test, the expected $U_{k,n,w}$ = 1.1300, the observed value is $U_{k,n,w}$ = 1.1245, and the p-value is $\text{p-value}_{U_{k,n,w}}$ = 0.7055. Again, the results do not indicate a significant difference between the distributions. In contrast, as shown in Fig.˜22, when running with PQMass using 10 regions (due to limited samples), the resulting $\chi_{\rm PQM}^{2}$ as well as the $\text{p-value}_{\rm PQM}$, confidently indicate the data does not come from the same underlying distribution.
@@ Appendix I Permutation tests
When PQMass is run for a single iteration (single tessellation) it is an exact statistical test, however the power of this test is considerably dependent on the particular arrangement of the Voronoi bins. To increase the power of the test, we repeat the test multiple times and verify that $\chi^{2}_{PQM}$ is $\chi^{2}$ distributed. In cases where data samples are limited, however, running multiple tests by ressampling and retessellating can be difficult. In those cases, in particular in some of our experiments, we can run many retessellations on the same sets of samples to maximize the use of existing samples and ensure we encounter especially discriminating configurations.
In this case, however, since the samples in the retessellations are the same, there is some shared information between tests and the $\chi^{2}_{PQM}$ values from these retessellations must no longer follow a $\chi^{2}$ distribution under the null hypothesis. This presents a challenge for understanding the meaning of a PQMass $\chi^{2}$ retessellation histogram. This issue can be easily eliminated via a permutation test (Good, 2013).
Any measurement on some dataset can be turned into an exact test on the null hypothesis that two samples are drawn from the same distribution using a permutation test. In brief, a permutation test involves taking some scalar measurement on two datasets, lets say ${\rm PQM}(x,y)$ where $x$ and $y$ are the two samples, and we re-run the measurement on a permutation of the datasets ${\rm PQM}(x^{\prime},y^{\prime})$ where $x^{\prime}$ and $y^{\prime}$ are a random permutation of the combined dataset $\{x,y\}$. The permuted samples are thus drawn (without replacement) from a mixture distribution of the two samples. Running many permutations one builds up a distribution of values for ${\rm PQM}(x^{\prime},y^{\prime})$ against which we may compare the original value ${\rm PQM}(x,y)$ to find a p-value. At an intuitive level, this test determines how unique the original arrangement of $x,y$ samples are under the measurement $\rm PQM$}.
The PQMass test is especially sensitive to the differences in two samples, making it an ideal choice for such a test. Specifically, we may now take PQMass run with many retessellations, take the mean of the reported $\chi^{2}$ values, and use this mean as our measurement on the two samples; now fully utilizing the discriminating power of the retessellations in an exact test. We perform a permutation test using PQMass in Fig.˜23. The histogram shows PQMass run on random permutations of the input samples (“permute stats” in the figure) while the vertical line shows the result for the input samples (“test stat” in the figure). In the null case, we achieve a p-value consistent with $U(0,1)$ while in the out-of-distribution case we reject the null with a low p-value.
In both cases, the mean of many re-retessellations on the original sample would have placed us well within the $\chi^{2}(99)$ distribution, showing how the permutation test is more powerful at rejecting the null, if such extra power is needed.
In practice, we find that, for high-dimensional problems, the retessellations without resampling very closely approximate a $\chi^{2}$ distribution (in the null case), though this is not the case in low dimensions. Since it is common in high-dimensional spaces for PQMass to be highly discriminating, it is often not necessary to perform the permutation test to reject the null; one can conservatively do so using the full $\chi^{2}$ distribution.
If, after a few retessellations, one finds that they cannot easily reject the null, then performing a permutation test can add more discrimination power at the cost of extra compute. Even with this extra computational cost, PQMass remains computationally feasible on large and high-dimensional datasets. We note again that with a permutation test it is possible to turn any metric on two samples into an exact two-sample test; for example, the energy distance may be used (Székely et al., 2004), but the computational complexity of that test is higher than PQMass.
@@
Figure 23: Permutation test for validity of taking the mean of many PQMass retessellations. Samples are drawn from a 100 dimension standard Gaussian, we use 1000 samples for each input. Shown are the $\chi^{2}(99)$ distribution (black dashed), mean of many retessellations (red vertical line), and the permutation test (purple histogram). Left: A null test which, as expected, fails to reject the null hypothesis. Right: An out of distribution test where one sample has been scaled by $1.$ The null hypothesis is indeed rejected.